The president of Sheffield’s Shoes forecasts the company will incur $1,560,000 of manufacturing overhead in the...

  1. The president of Sheffield’s Shoes forecasts the company will incur $1,560,000 of manufacturing overhead in the coming year. The company also expects the following results for its operations in the coming year:

Direct labor hours 192,000

Direct labor cost $ 2,400,000

Machine hours 60,000

Compute overhead allocation rates based on the following bases:

Direct labor hours

Direct labor dollars

Machine hours

If the company decides to use direct labor hours as its basis for overhead allocation, how

much overhead would be allocated to a product that requires 5,000 hours of direct labor?
Answer #1

Predetermined overhead rate = Total estimated overhead costs / total estimated quantity of the overhead allocation base

Predetermined overhead rate = $1,560,000/192000 = $8.13 per Direct labor hour

Predetermined overhead rate = $1,560,000/2,400,000 = $0.65 per Direct labor cost

Predetermined overhead rate = $1,560,000/60,000 = $26 per Machine hour

Overhead allocated to the product = direct labor hours incurred * Predetermined overhead rate per Direct labor hour

= 5000*$8.13

= $40,650